# Exploration of morphological coherence in open clusters with “core-shell” structure

### v1

@@ Abstract
Context. The morphology of open clusters plays a major role in the study of their dynamic evolution. The study of their morphological coherence (a 3D difference between the inner and outer morphologies of open clusters) allows for a better understanding of the morphological evolution of open clusters.
@@
Methods. We employ the ellipsoid fitting method to delineate the 3D spatial structure of the sample clusters while using the morphological dislocation (MD) defined in our previous work and the ellipticity ratio (ER) of the clusters’ inner and outer structures to characterize the morphological coherence of the sample clusters.
Results. The results show an inverse correlation between the ER of the sample clusters and the number of their members, indicating that sample clusters with much more elliptical external morphology than internal shape have generally a large number of members. Meanwhile, a slight shrinking of the MD of the sample clusters with their members’ number may shed light on the significant role of the gravitational binding of the sample clusters in maintaining their morphological stability. Moreover, there are no correlations between the MD and ER of the sample clusters and their age. They are also not significantly correlated with the X-axis, the Y-axis, their orbital eccentricities, and the radial and vertical forces on them. However, the ER of the sample clusters displays some fluctuations in the distributions between it and the above covariates, implying that the morphologies of the sample clusters are sensitive to the external environment if sample effects are not taken into account. Finally, the analysis of the 3D spatial shapes of sample clusters with a small ER or a large ER demonstrates that the number of members lays an important foundation for forming a dense internal system for sample clusters. At the same time, the MD of the sample clusters can serve well as an indicator of their morphological stability, which is built on a certain amount of member stars.
Conclusions. We present a new insight into the morphological coherence of open clusters, which would be attributed to the combination of their gravitational binding capacity and external environmental perturbations.
@@ 1 Introduction
Open clusters are an important part of the Milky Way, mainly distributed on the Galactic plane. As a consequence, it is of great significance for the study of the Milky Way, such as in tracing the structures of the galactic spiral arms (Dias & Lépine, 2005). In the evolution of the open clusters, the number of their members must be changed due to star evolution (Portegies Zwart et al., 2010), collisional dynamics, and tidal perturbations (Krause et al., 2020). Their shapes can visually indicate this change. In addition, the morphology of the open clusters may inherit morphological imprints of their formation period (Krause et al., 2020), which may help us to follow their formation mechanism. Therefore, the morphology of open clusters plays a vital role in the formation and evolution of open clusters.
@@
The typical shape presented in the clusters is thought to be the core-corona structure (e.g., Artyukhina & Kholopov, 1966; Kholopov, 1969; Chen et al., 2004; Zhai et al., 2017), which was named as the layered structure in our study (Hu et al., 2023). Based on this structure, in our series of papers (Hu et al., 2021a, b, 2023), we defined for the first time a parameter named morphological dislocation of open clusters, to assess the stability of the two-dimensional morphology of the open clusters in the statistical study of their morphological laws and properties.
@@
As long as open clusters can survive in their embedded phase (Lada & Lada, 2003), the physical processes that may change the morphology of the open clusters are roughly as follows. As a cluster evolves, the distribution of stellar members in its core is likely modified by star evolution, mass segregation (Vesperini et al., 2009; Portegies Zwart et al., 2010; Evans & Oh, 2022; Noormohammadi et al., 2023), and internal gravitational interaction among the members. Subsequently, stellar evaporation (Portegies Zwart et al., 2010; Reina-Campos et al., 2023), expansion resulted from dynamical evolution (Krause et al., 2020), possible intrusion of non-native stars (Hu et al., 2022), and external perturbations (Spitzer, 1958; Portegies Zwart et al., 2010), e.g., Galactic tidal stripping due to the Galactic potential (Heggie & Hut, 2003; Ernst et al., 2015; McDivitt et al., 2022; Angelo et al., 2023), Galactic differential rotation, and gravitational perturbations from giant molecular clouds, would alter the spatial structure of its outer region. In the process above, the internal and external morphologies of the open cluster will gradually show differences. Will such differences have an impact on cluster evolution and cluster morphological stability?
Although many efforts have been made to study the morphology of open clusters (e.g., Jeans, 1916; Artyukhina & Kholopov, 1966; Kholopov, 1969; Oort, 1979; Pandey et al., 1990; Bergond et al., 2001; Nilakshi et al., 2002; Chen et al., 2004; Cartwright & Whitworth, 2004; Santos et al., 2005; Kharchenko et al., 2009; Zhai et al., 2017; Dib et al., 2018; Hetem & Gregorio-Hetem, 2019; Pang et al., 2021; Tarricq et al., 2022; Zhong et al., 2022), few have studied the morphological issue mentioned due to the lack of parameters to assess it. This has to some extent limited our comprehensive and in-depth understanding and study of the open clusters.
@@
In order to find out the cluster morphology question, we define for the first time in this study the parameter of the inner and outer differences in cluster morphology to explore the question. That is “morphological coherence” (a 3D difference between the morphology of the clusters’ core region and their external morphology, similar to the morphological dislocation of Hu et al. (2021b)), a new parameter of cluster morphology, which can indicate clusters’ morphological stability. The morphological coherence of the open clusters can be considered as the morphological stability for individual clusters. In general, the poor morphological stability of the open clusters corresponds to poor morphological coherence, which can be shown visually in the distribution of their member stars. Furthermore, this morphological coherence is constantly changing as the clusters evolve. Therefore, the morphological coherence of the open clusters can lay an important foundation for understanding their evolutionary processes.
@@
In the present work, we aim to explore the morphological coherence of the open clusters with core-shell structure in three-dimensional space up to 1 kpc from the Sun, based on cluster catalogs in the literature. By fitting a 3D ellipsoid to the sample clusters to obtain their fitting parameters and then calculate the morphological coherence, by which we attempt to investigate the morphological coherence of open clusters in terms of their basic parameters (age and the number of member stars), spatial positions, and an external environment. This will further enrich the morphological study and also help to provide observational evidence for tracing the formation and evolution mechanisms of open clusters. In the following, we describe the data and methodology in Sect. 2. Some of the results we obtained are presented in Sect. 3. Finally, Sect. 4 summarizes our work and lists some important conclusions.
@@ 2 Data and method
On the ground of the Gaia Data Release 3 (DR3) data, Hunt & Reffert (2023) search blindly for and identify member stars of open clusters within all-sky based on the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithms. They searched 7167 clusters, which included 2387 new candidates and 4105 highly reliable clusters in their catalog, and provided basic parameters and membership lists for all clusters. The sample in this work was selected from their clusters catalog. We filtered 682 sample clusters from the 7167 clusters by six constraints: 1) “kind” $=$ “o”; 2) S/Ns $\geq$ 5; 3) class_50 (median CMD classifications) $\geq$ 0.5; 4) membership probability $\geq$ 0.5; 5) parallax ($\varpi$$\geq$ 1; 6) members’ number (N) $\geq$ 50. The first three terms can ensure the objects we elected are highly reliable open clusters (Hunt & Reffert, 2023), and the last three conditions can meet the requirements of our sample. It should be noted that the term of member probability $\geq$ 0.5 is able to remove low-quality members (Hunt & Reffert, 2023), which can guarantee the reliability of member stars. The cut can also remove stars in tidal tails (Hunt & Reffert, 2023), but not all, so some of our sample clusters still include gravitationally bound and non-gravitational bound member stars. Finally, a three-dimensional ellipsoidal fit was used to filter the sample. The cut criterion is that the errors of the three axes of the fitted ellipsoid for any sample cluster, both core and shell, must be smaller than the length of the three axes, respectively. This is because it can guarantee the sample clusters we elected, both core and shell, have stable ellipsoid structures. If the error of any axis of clusters fitted is more than the length of their axis, the fitting ellipsoid structures of the clusters may be unstable or fake. Moreover, the fits were checked visually, frame by frame, to exclude pseudo-ellipsoid fits. These pseudo-ellipsoid fits meet our filter condition, but the member stars of clusters fitted by them is almost exclusively distributed around one patch of the surface of the fitted ellipsoids, especially the ellipsoid fitted of the shell, rather than distributed around the entire surface of the fitted ellipsoids. In this way, we obtained a total of 132 sample clusters, all of which have a “core-shell” structure (shown in Fig. 1) and fit the ellipsoidal model (displayed in Fig. 2), see more details in the next subsection. These 132 open clusters are the research subject of this work.
Figure 1: The spatial distribution of the member stars of the cluster (Collinder 135) in the heliocentric Cartesian coordinate system. The red open circles denote the members within the core region of the cluster, with the blue open circles being the stars in its outer region.
@@ 2.1 Ellipsoid fitting for the core-shell structure of sample clusters
To fit an ellipsoid to the core-shell structure of sample clusters, we need to first determine the three-dimensional spatial distribution of the clusters’ members and then delineate their core-shell structure. Before obtaining the space distribution, we calculated the heliocentric Cartesian coordinates (X, Y, Z) for each star member of the sample clusters. This calculation was performed with the Python Astropy package (Astropy Collaboration et al., 2013, 2018).
@@
In light of the core-shell structure of star clusters and the layered structure of open clusters detected by Hu et al. (2023), we divided the stellar members of each sample cluster into two different regions (including a core and outer regions, i.e., the core represents those stars within the radius of the half number, with the outer field indicating those members out of this radius) based on the radius of the half number of the cluster member stars, shown in Fig. 1. The starting point of this half-number radius is the centroid of the 3D spatial distribution of the member stars, which is obtained from the three-dimensional nuclear density assessment after the members’ distance correction. We set out to delineate the morphologies of the two different regions of each sample cluster in 3D space by ellipsoid fitting. The fitting ellipsoids of the sample clusters thus were obtained, as shown in Fig. 2 and 3.
@@ 2.2 Characterization of the morphological coherence of the sample clusters
The morphological coherence of the sample clusters, in fact, refers to a 3D difference between the inner and outer morphologies in this work. It is essentially determined by their core-shell structures, mainly in the 3D spatial distributions of member stars in their inner and outer regions. Any variable that could measure this disparity between the inner and outer regions can indicate the cluster’s morphological coherence. The morphological dislocation is probably an appropriate indicator. The morphological dislocation of open clusters (hereafter MD), i.e., the offset between the centers of the core and outer regions, is important in the study of their layered structure and morphological stability. Its definition was originally taken from our previous paper of this series (Hu et al., 2021b), which was performed in 2D space. In this work, we employ it to probe the morphological coherence of the clusters. Unlike previous work, it is operated in three dimensions in the current work. To clearly describe this new definition, the detailed equation is listed below:
$d=\frac{\sqrt{(X_{outer}-X_{core})^{2}+(Y_{outer}-Y_{core})^{2}+(Z_{outer}-Z_{core})^{2}}}{a_{outer}}$ (1)
@@
$d$ is a dimensionless physical quantity denoting the morphological dislocation of the sample clusters in 3D space. The $d$ specifically refers to the system centers’ relative offset between the core and the outer regions of the clusters. $X_{core}$, $Y_{core}$ and $Z_{core}$ denote the center coordinates of the fitting ellipsoid in the core regions of the clusters (see the solid green curves in Fig. 2 and 3), while $X_{outer}$, $Y_{outer}$ and $Z_{outer}$ refer to the center coordinates of the fitting ellipsoid in their outer regions, see the solid purple curves in Fig. 2 and 3. $a_{outer}$ represents the semi-major axis of the fitted ellipsoid of their outer fields. The error of $d$ is calculated based on error propagation. The left panel of Fig. 4 shows the distribution between $d$ (MD) and its error ($\sigma_{d}$).
@@
To fully explore the morphological coherence within the sample clusters, we also took the ellipticity ratio ($e_{core}$/$e_{outer}$) of their core-shell structures (hereafter ER) to describe the morphological difference between their inner and outer structures, characterizing the morphological coherence of the clusters. The detailed ellipticity formula is as follows:
$\qquad\qquad\qquad\qquad e_{core}=1-\frac{c_{core}}{a_{core}}$ (2)
$\qquad\qquad\qquad\qquad e_{outer}=1-\frac{c_{outer}}{a_{outer}}$ (3)
@@
$e_{core}$ and $e_{outer}$ refer to the ellipticities of the core and shell regions of the sample clusters, respectively. $a_{core}$ and $c_{core}$ represent the semi-major and semi-minor axes of the fitted ellipsoids of their core fields, respectively, while $a_{outer}$ and $c_{outer}$ correspond to the semi-major and semi-minor axes of the fitted ellipsoids of their outer regions, respectively. The errors of $e_{core}$, $e_{outer}$, and $e_{core}$/$e_{outer}$ are also calculated according to error transmission. The right panel of Fig. 4 displays the distribution between $e_{core}$/$e_{outer}$ (ER) and its error ($\sigma_{e_{core}/e_{outer}}$).
@@
Whether the ER ($e_{core}$/$e_{outer}$) is greater than 1 or less than 1, the inner and outer morphologies of the sample clusters are different. It is noted that the ER greater than 1 indicates that the core of the clusters is more elliptical than the shell, and conversely, the ER less than 1 suggests that the core of the sample clusters is less elliptical compared to their shell. Angelo et al. (2023) pointed out that clusters tend to form a compact core along their dynamical evolution process, which may be a result of the movement of massive members toward the clusters’ core based on two-body interactions. Meanwhile, the sphericity of the cores also increases (Chen et al., 2004). The sample clusters with ER less than 1 or greater than 1 might face two scenarios. Clusters with more elliptical external morphology than internal morphology likely have undergone a two-body relaxation process (Portegies Zwart et al., 2010), tending to have a circular internal structure (Heggie & Hut, 2003; Chen et al., 2004) and an advanced evolution status (Angelo et al., 2023). In addition, this shape presents likely a large Roche volume filling factor (Angelo et al., 2023) representing the degree of tidal filling of a cluster (Santos et al., 2020; Angelo et al., 2023), which means that the clusters with this morphology are subject to both external tidal fields and internal evolution. On the contrary, the sample clusters with an elliptical core and a circular shell may be suffering from an internal two-body relaxation.

### v2

# Exploration of morphological coherence in open clusters with a “core-shell” structure

@@ Abstract
Context. The morphology of open clusters plays a major role in the study of their dynamic evolution. The study of their morphological coherence, namely, the three-dimensional (3D) difference between the inner and outer morphologies of open clusters, allows us to obtain a better understanding of the morphological evolution of open clusters.
@@
Methods. We employed the ellipsoid fitting method to delineate the 3D spatial structure of the sample clusters, while using the morphological dislocation (MD) defined in our previous work and the ellipticity ratio (ER) of the clusters’ inner and outer structures to characterize the morphological coherence of the sample clusters.
Results. The results show an inverse correlation between the ER of the sample clusters and the number of their members, indicating that sample clusters with a much more elliptical external morphology than internal shape generally tend to host a large number of members. Meanwhile, a slight shrinking of the MD of the sample clusters with their members’ number may shed light on the significant role of the gravitational binding of the sample clusters in maintaining their morphological stability. Moreover, there are no correlations between the MD and ER of the sample clusters and their age. They are also not significantly correlated with the X-axis, the Y-axis, their orbital eccentricities, and the radial and vertical forces on them. However, the ER of the sample clusters displays some fluctuations in the distributions between it and the above covariates, implying that the morphologies of the sample clusters are sensitive to the external environment if sample effects are not taken into account. Finally, the analysis of the 3D spatial shapes of sample clusters with a small ER or a large ER demonstrates that the number of members lays an important foundation for forming a dense internal system for sample clusters. At the same time, the MD of the sample clusters can serve well as an indicator of their morphological stability, which is built upon a certain amount of member stars.
Conclusions. We present a new insight into the morphological coherence of open clusters, attributed to the combination of their gravitational binding capacity and external environmental perturbations.
@@ 1 Introduction
Open clusters, which are mainly distributed along the Galactic plane, make up an important part of the Milky Way. As a consequence, it is of great significance for the study of the Milky Way, such as tracing the structures of the galactic spiral arms (Dias & Lépine, 2005). In the evolution of the open clusters, the number of their members must change due to star evolution (Portegies Zwart et al., 2010), collisional dynamics, and tidal perturbations (Krause et al., 2020). Their shapes can visually indicate this change. In addition, the morphology of the open clusters may inherit morphological imprints of their formation period (Krause et al., 2020), which can allow us to follow their formation mechanism. Therefore, the morphology of open clusters plays a vital role in the formation and evolution of open clusters.
@@
The typical shape presented in the clusters is thought to be the core-corona structure (e.g., Artyukhina & Kholopov, 1966; Kholopov, 1969; Chen et al., 2004; Zhai et al., 2017), referred to as the layered structure in our previous study (Hu et al., 2023). Based on this structure, in our series of papers (Hu et al., 2021a, b, 2023), we defined, for the first time, a parameter denoted as the “morphological dislocation” of open clusters to assess the stability of the two-dimensional (2D) morphology of the open clusters in the statistical study of their morphological laws and properties.
@@
As long as open clusters can survive in their embedded phase (Lada & Lada, 2003), the physical processes that may change the morphology of the open clusters can be roughly described as follows. As a cluster evolves, the distribution of stellar members in its core is likely modified by star evolution, mass segregation (Vesperini et al., 2009; Portegies Zwart et al., 2010; Evans & Oh, 2022; Noormohammadi et al., 2023), and internal gravitational interaction among its members. Subsequently, stellar evaporation (Portegies Zwart et al., 2010; Reina-Campos et al., 2023), expansion resulted from dynamical evolution (Krause et al., 2020), possible intrusion of non-native stars (Hu et al., 2022), and external perturbations (Spitzer, 1958; Portegies Zwart et al., 2010), for instance, Galactic tidal stripping due to the Galactic potential (Heggie & Hut, 2003; Ernst et al., 2015; McDivitt et al., 2022; Angelo et al., 2023), Galactic differential rotation along with, gravitational perturbations from giant molecular clouds, would alter the spatial structure of its outer region. In the process above, the internal and external morphologies of the open cluster will gradually show differences. In this work we consider whether such differences have an impact on cluster evolution and cluster morphological stability. Numerous efforts have been undertaken to study the morphology of open clusters (e.g., Jeans, 1916; Artyukhina & Kholopov, 1966; Kholopov, 1969; Oort, 1979; Pandey et al., 1990; Bergond et al., 2001; Nilakshi et al., 2002; Chen et al., 2004; Cartwright & Whitworth, 2004; Santos et al., 2005; Kharchenko et al., 2009; Zhai et al., 2017; Dib et al., 2018; Hetem & Gregorio-Hetem, 2019; Pang et al., 2021; Tarricq et al., 2022; Zhong et al., 2022); however, very few have studied the morphological issue mentioned above due to the lack of parameters to assess it. This has limited (to some extent) our comprehensive and in-depth understanding and study of the open clusters.
@@
To address the cluster morphology question, here we define, for the first time, the parameter of the inner and outer differences in cluster morphology. This “morphological coherence”, namely, the 3D difference between the morphology of the clusters’ core region and their external morphology, similar to the morphological dislocation of Hu et al. (2021b)) is a new parameter of cluster morphology that can indicate clusters’ morphological stability. The morphological coherence of the open clusters can be considered as the morphological stability for individual clusters. In general, the poor morphological stability of the open clusters corresponds to poor morphological coherence, which can be shown visually in the distribution of their member stars. Furthermore, this morphological coherence is constantly changing as the clusters evolve. Therefore, the morphological coherence of the open clusters can lay an important foundation for understanding their evolutionary processes.
@@
In the present work, we aim to explore the morphological coherence of the open clusters with core-shell structure in 3D space up to 1 kpc from the Sun, based on cluster catalogs from the literature. By fitting a 3D ellipsoid to the sample clusters to obtain their fitting parameters and then calculate the morphological coherence, we attempt to investigate the morphological coherence of open clusters in terms of their basic parameters (age and the number of member stars), spatial positions, and external environment. This will further enrich our morphological study and also help to provide observational evidence for tracing the formation and evolution mechanisms of open clusters. In the following, we describe the data and methodology in Sect. 2. We present a summary of our results in Sect. 3. Finally, Sect. 4 summarizes our work and gives some important conclusions.
@@ 2 Data and method
On the basis of the Gaia Data Release 3 (DR3) data, Hunt & Reffert (2023) carried out a blind search for member stars of open clusters within all-sky based on the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithms, with an aim to identify them. They searched 7167 clusters, which included 2387 new candidates and 4105 highly reliable clusters in their catalog, and provided basic parameters and membership lists for all clusters. The sample in this work was selected from their clusters catalog. We filtered 682 sample clusters from the 7167 clusters by six constraints: 1) “kind” $=$ “o”; 2) S/Ns $\geq$ 5; 3) class_50 (median CMD classifications) $\geq$ 0.5; 4) membership probability $\geq$ 0.5; 5) parallax ($\varpi$$\geq$ 1; 6) members’ number (N) $\geq$ 50. The first three terms can ensure the objects we elected are highly reliable open clusters (Hunt & Reffert, 2023), and the last three conditions can meet the requirements of our sample. It should be noted that the term of member probability $\geq$ 0.5 is able to remove low-quality members (Hunt & Reffert, 2023), which can guarantee the reliability of member stars. The cut may also remove stars in tidal tails (Hunt & Reffert, 2023), but not all, so some of our sample clusters still include gravitationally bound and non-gravitational bound member stars. Finally, a 3D ellipsoidal fit was used to filter the sample. The cut criterion is that the errors of the three axes of the fitted ellipsoid for any sample cluster, both core and shell, must be smaller than the length of the three axes, respectively. This guarantees the sample clusters we elected, both core and shell, have stable ellipsoid structures. If the error of any axis of clusters fitted is more than the length of their axis, the fitting ellipsoid structures of the clusters may be unstable or fake. Moreover, the fits were checked visually, frame by frame, to exclude pseudo-ellipsoid fits. These pseudo-ellipsoid fits meet our filter condition, but the member stars of clusters fitted by them is almost exclusively distributed around one patch of the surface of the fitted ellipsoids (especially the ellipsoid fitted of the shell), rather than distributed around the entire surface of the fitted ellipsoids. In this way, we obtained a total of 132 sample clusters, all of which have a “core-shell” structure (shown in Fig. 1) and fit the ellipsoidal model (displayed in Fig. 2, with more details given in the next subsection). These 132 open clusters are the research subject of this work.
Figure 1: Spatial distribution of the member stars of the cluster (Collinder 135) in the heliocentric Cartesian coordinate system. The red open circles denote the members within the core region of the cluster, with the blue open circles being the stars in its outer region.
@@ 2.1 Ellipsoid fitting for the core-shell structure of sample clusters
To fit an ellipsoid to the core-shell structure of sample clusters, we need to first determine the 3D spatial distribution of the clusters’ members and then delineate their core-shell structure. Before obtaining the space distribution, we calculated the heliocentric Cartesian coordinates (X, Y, Z) for each star member of the sample clusters. This calculation was performed with the Python Astropy package (Astropy Collaboration et al., 2013, 2018).
@@
In light of the core-shell structure of star clusters and the layered structure of open clusters detected by Hu et al. (2023), we divided the stellar members of each sample cluster into two different regions (including a core and outer regions, i.e., the core represents those stars within the radius of the half number, with the outer field indicating those members out of this radius) based on the radius of the half number of the cluster member stars (shown in Fig. 1). The starting point of this half-number radius is the centroid of the 3D spatial distribution of the member stars, which is obtained from the 3D nuclear density assessment after the members’ distance correction. We set out to delineate the morphologies of the two different regions of each sample cluster in 3D space by ellipsoid fitting. The fitting ellipsoids of the sample clusters thus were obtained, as shown in Fig. 2 and 3.
@@ 2.2 Characterization of the morphological coherence of the sample clusters
The morphological coherence of the sample clusters, in fact, refers to a 3D difference between the inner and outer morphologies in this work. It is essentially determined by their core-shell structures, mainly in the 3D spatial distributions of member stars in their inner and outer regions. Any variable that could measure this disparity between the inner and outer regions can indicate the cluster’s morphological coherence. The morphological dislocation is probably an appropriate indicator. The morphological dislocation of open clusters (hereafter MD), namely, the offset between the centers of the core and outer regions, is important in the study of their layered structure and morphological stability. Its definition was originally taken from our previous paper of this series (Hu et al., 2021b), which was performed in 2D space. In this work, we employ it to probe the morphological coherence of the clusters. Unlike our previous work, here it was operated in three dimensions. To clearly describe this new definition, the detailed equation is listed below:
$d=\frac{\sqrt{(X_{outer}-X_{core})^{2}+(Y_{outer}-Y_{core})^{2}+(Z_{outer}-Z_{core})^{2}}}{a_{outer}}.$ (1)
@@
Here, $d$ is a dimensionless physical quantity denoting the morphological dislocation of the sample clusters in 3D space. The value of $d$ specifically refers to the system centers’ relative offset between the core and the outer regions of the clusters. Then, $X_{core}$, $Y_{core}$ and $Z_{core}$ denote the center coordinates of the fitting ellipsoid in the core regions of the clusters (see the solid green curves in Figs. 2 and 3), while $X_{outer}$, $Y_{outer}$ and $Z_{outer}$ refer to the center coordinates of the fitting ellipsoid in their outer regions (see the solid purple curves in Figs. 2 and 3). $a_{outer}$ represents the semi-major axis of the fitted ellipsoid of their outer fields. The error of $d$ is calculated based on error propagation. The left panel of Fig. 4 shows the distribution between $d$ (MD) and its error ($\sigma_{d}$).
@@
To fully explore the morphological coherence within the sample clusters, we also took the ellipticity ratio ($e_{core}$/$e_{outer}$) of their core-shell structures (hereafter, ER) to describe the morphological difference between their inner and outer structures, characterizing the morphological coherence of the clusters. The detailed ellipticity formula is as follows:
$\qquad\qquad\qquad\qquad e_{core}=1-\frac{c_{core}}{a_{core}},$ (2)
$\qquad\qquad\qquad\qquad e_{outer}=1-\frac{c_{outer}}{a_{outer}},$ (3)
@@
Here, $e_{core}$ and $e_{outer}$ refer to the ellipticities of the core and shell regions of the sample clusters, respectively; $a_{core}$ and $c_{core}$ represent the semi-major and semi-minor axes of the fitted ellipsoids of their core fields, respectively, while $a_{outer}$ and $c_{outer}$ correspond to the semi-major and semi-minor axes of the fitted ellipsoids of their outer regions, respectively. The errors of $e_{core}$, $e_{outer}$, and $e_{core}$/$e_{outer}$ are also calculated according to error transmission. The right panel of Fig. 4 displays the distribution between $e_{core}$/$e_{outer}$ (ER) and its error ($\sigma_{e_{core}/e_{outer}}$).
@@
When the ER ($e_{core}$/$e_{outer}$) is greater than 1 or less than 1, the inner and outer morphologies of the sample clusters are different. It is noted that the ER greater than 1 indicates that the core of the clusters is more elliptical than the shell and, conversely, the ER less than 1 suggests that the core of the sample clusters is less elliptical compared to their shell. Angelo et al. (2023) pointed out that clusters tend to form a compact core along their dynamical evolution process, which may be a result of the movement of massive members toward the clusters’ core based on two-body interactions. Meanwhile, the sphericity of the cores also increases (Chen et al., 2004). The sample clusters with ER less than 1 or greater than 1 might face two scenarios. Clusters with more elliptical external morphology than internal morphology likely have undergone a two-body relaxation process (Portegies Zwart et al., 2010), tending to have a circular internal structure (Heggie & Hut, 2003; Chen et al., 2004) and an advanced evolution status (Angelo et al., 2023). In addition, this shape presents likely a large Roche volume filling factor (Angelo et al., 2023) representing the degree of tidal filling of a cluster (Santos et al., 2020; Angelo et al., 2023), which means that the clusters with this morphology are subject to both external tidal fields and internal evolution. On the contrary, the sample clusters with an elliptical core and a circular shell may be suffering from an internal two-body relaxation.